A diploid species has a total of 40 chromosomes. How many chromosomes would be found in...

  1. A diploid species has a total of 40 chromosomes. How many chromosomes would be found in a monosomy, trisomy, and autotriploid?

    51; 78; 53

    39; 41; 60

    51; 53; 104

    41; 39; 60

    51; 53; 78
Answer #1

Monosomy: from homologous chromosome pair only one chromosome is present.

2n-1= 40-1=39 chromosome

Trisomy: in homologous chromosome pair, addition chromosome is present

2n+1=40+1=41

Autotriploidy: three chromosome set are present

3*n=3*20=60

Answer: 39,41,60